OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4 1 −1 −5 −25

OpenStudy (jhannybean):

work this problem the same way as the last one!:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got -9 i did it wrong

OpenStudy (jhannybean):

3x + 3 = -2x +2 -4 Add +2x to both sides 3x + 2x + 3 = 2 - 4 5x + 3 = -2 subtract -3 from both sides 5x + 3 - 3 = -2 - 3 Simplify 5x = -5 divide LHS and RHS by 5 5x/5 = -5/5 x = -1

OpenStudy (jhannybean):

ty :)